# Investigation of the effects of superhydrophobic surface treatment on the dynamics of the flow in the near wake of a sphere using spatial dynamic mode decomposition

### v1

@@ 5 Concluding Remarks
The streamwise and transverse velocities in a plane in the near wake of spheres with and without superhydrophobic surface treatment were measured using 2C-2D MCCD-PIV, with a large number of instantaneous velocity fields. The mean planar velocities were then determined and used to calculate the instantaneous velocity fluctuations, which were decomposed into DMD modes by applying dynamic mode decomposition in the streamwise direction. The resulting DMD modes were sorted by decay rate, from slowest to fastest, and the first three DMD modes for each sphere were compared.
The first two DMD modes in the REF and PRS cases were found to be qualitatively and quantitatively similar, with matching DMD mode shapes, decay rates, and wavelengths. The third DMD modes in these cases were qualitatively similar, albeit with the PRS case being somewhat asymmetric relative to the REF case. This asymmetry is likely due to the spherical distribution of the pores on the sphere resulting in slightly different flow conditions on either side of the sphere in the measured plane. While the first mode of the SHS sphere was qualitatively similar to those of the REF and PRS cases, the wavelength was significantly shorter. The second and third DMD modes of this case were radically different from the untreated spheres, suggesting a significant influence from the plastron behaviour on the dynamics in the near wake, greater three-dimensionality in the wake of the SHS sphere than in the REF and PRS cases.
The current work shows that the addition of superhydrophobic surface treatment has a significant effect on the dynamics of the flow in the near wake of a sphere, which is consistent with the variation in drag effects reported in previous studies. Thus, the drag effects of superhydrophobic surface treatment in turbulent flow are not just a function of the partial slip generated at the surface when the plastron is maintained, but also of the effect that the plastron has on the flow as it is deformed by the velocity and pressure fluctuations.
@@ Acknowledgements
Shaun Davey gratefully acknowledges the support of the Australian Commonwealth Government through a Research Training Program (RTP) Scholarship.
This research was undertaken with the support of resources provided via a Monash HPC-NCI Merit Allocation from the National Computational Infrastructure (NCI Australia), an NCRIS-enabled capability supported by the Australian Government.

### v2

@@ 5 Concluding Remarks
The streamwise and transverse velocities in the near wake of spheres with and without superhydrophobic surface treatment were measured using 2C-2D MCCD-PIV. The mean planar velocities were then determined and used to calculate the instantaneous velocity fluctuations, which were separated into domains above and below the sphere’s centreline, centred on the shear layer. DMD was performed on the fluctuations in each domain, using a shift in the streamwise direction as an analogue for a temporal separation. The DMD modes in each domain were compared and adjusted for phase shifts before averaging. The resulting DMD modes were sorted by energy share, with long wavelengths assigned lower rankings. The complex cross-correlation was used to identify equivalent modes across spheres to assess the effect of pores and superhydrophobic surface treatment on the shear-layer dynamics. The leading four DMD modes of each sphere represent equivalent instabilities, with the wavelengths, decay rates and mode shapes altered by the superhydrophobic surface treatment. In addition to these modes, the untreated spheres featured similar fifth and sixth modes, with the sixth mode having a wavelength significantly longer than the measurement domain despite having the highest energy share. Similarly, the fifth and sixth modes of the SHS sphere had wavelengths significantly longer than the measurement domain, with long structures featuring spatially regular increases in intensity. These smaller structures may be the result of periodic changes in the decaying turbulence due to air dissolving into the fluid as the plastron dissipates. Where the leading POD modes of the SHS sphere were contaminated by the increased three‑dimensionality of the dominant untreated‑sphere structures, the current application of DMD, using spatial advection as an analogue for time, was able to identify the primary instabilities in the wake’s shear layer. While the addition of pores to the sphere’s surface had some effect on these instabilities, the superhydrophobic surface treatment produced more substantial changes. Additionally, while all three spheres included long-wavelength modes with high energy shares, the SHS sphere featured small, regularly spaced structures within these modes.
@@ Acknowledgements
The authors acknowledge the support of the Australian Research Council (ARC) in funding this research through Linkage Infrastructure, Equipment and Facilities (LIEF) and Discovery grants. Shaun Davey gratefully acknowledges the support of the Australian Commonwealth Government through a Research Training Program (RTP) Scholarship. This research was undertaken with the support of resources provided via a Monash HPC-NCI Merit Allocation from the National Computational Infrastructure (NCI Australia), an NCRIS-enabled capability supported by the Australian Government.